OpenStudy (anonymous):

Part 1: Find the solution to the equation log16 32 = x + 5. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (anonymous):

log16(32) = log 2^4 (2^5) = 5/4* log 2(2) = 5/4. The term outside the bracket is the base of the log. 5/4= x +5 , x = -5 + 5/4 = -15/4